OpenStudy (anonymous):

10L of a gas collected over 1000ml of water at 300K after collection weight of the water decreased by 0.24g The vapour pressure of water is?

mathslover (mathslover):

wht abt me ?

mathslover (mathslover):

so yahoo ! it is your home work right ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

sorry i forgot u @mathslover

mathslover (mathslover):

np

OpenStudy (anonymous):

Can u help me

OpenStudy (anonymous):

@asnaseer

OpenStudy (anonymous):

@Tushara

OpenStudy (anonymous):

no wat happens..............partial pressure of liquid = partial pressure of gas?

OpenStudy (anonymous):

the answer i got is way too small to be right, wat is the answer?

OpenStudy (anonymous):

@Tushara the asnwer should be 7

OpenStudy (anonymous):

oh ok, sorry, m not getting the right answer, cant help you in dis one

OpenStudy (anonymous):

@AravindG

OpenStudy (aravindg):

vapour pressure =molecular weight/2

OpenStudy (anonymous):

so we need MM

OpenStudy (aravindg):

yep try to use the data given

OpenStudy (aravindg):

do u hav final answeR?

OpenStudy (anonymous):

yup the asnwer should be 7

OpenStudy (aravindg):

wait a sec i ,ade a mistake

OpenStudy (anonymous):

i think u shud do it this way............. for water vapour pv=(w/m)rt u van get v using p1v2=p2v2 for that gas

OpenStudy (aravindg):

vapour density =molecular weight/2

OpenStudy (aravindg):

srry i read the qn wrong

OpenStudy (aravindg):

seems @A_Avinash_Goutham's working is good

OpenStudy (anonymous):

wait u wont get volume that way......u dont have initial pressure :P

OpenStudy (anonymous):

@heena

OpenStudy (anonymous):

@chmvijay

OpenStudy (anonymous):

hey u have to find vapour pressure so do only vapour presure=molecular weight of water/2 =18/2 =6

OpenStudy (anonymous):

as formular of water dosent change in any state -_-

OpenStudy (anonymous):

18/2=9

OpenStudy (anonymous):

@heena

OpenStudy (anonymous):

sorry yea 9 ^.^

OpenStudy (anonymous):

so next

OpenStudy (anonymous):

nooooooooooooooooo u never do qn by urself :/ now allong with ur qn show us ur attempt also -_-

OpenStudy (anonymous):

i want u to tell wat should i do i will solve it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then plzz tell

OpenStudy (anonymous):

@heena

OpenStudy (anonymous):

wat ? i guess for this we already shown u formula ...

OpenStudy (anonymous):

what means for the same qn or other qn u are going to give us?

OpenStudy (kayne):

@heena can you please send me a link on the formula you've used above - (molecular weight/2) for vapour pressure? I've been looking for it on the net but I can't get a grip on it. Thanks :)